OpenStudy (anonymous):

Find the polynomial function with roots 11 and 2i.

OpenStudy (jdoe0001):

for a polynomial of say \( x^2+2x+1 \) what would you say are its roots?

OpenStudy (jdoe0001):

do you even know what a 'root' of an expression is? :|

OpenStudy (anonymous):

2 and 1 I think

OpenStudy (jdoe0001):

well, 1 :|, because \( x^2+2x+1 \) = (x+1)(x+1)=0, x=1 and x=1, so it has 2 roots, both are 1 in this case

OpenStudy (jdoe0001):

a root of a polynomial means, if you equate it to 0, that is \( x^2+2x+1 =y \implies x^2+2x+1=0\), that is, the function is touching the X-axis, what values does X have

OpenStudy (jdoe0001):

you usually end up with one or more binomials, like the one above "x+1)(x+1)=0", so in this case your function where 'y=0', you'd have 2 binomials, like $$ \boldsymbol{i=\sqrt{-1}}\\ \color{blue}{(x-11)(x-2\sqrt{-1})=0}\\ x=11, x=2\sqrt{-1} = 2i $$

OpenStudy (jdoe0001):

if you multiply/expand those 2 binomials, you'd get your equation :)